# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 5.1. Tight non-fillable structures on spheres
We first construct tight, non-fillable contact structures on $\mathbb{S}^{2n+1}$, for $n\geq 2$ to prove Theorem B.
@@
We consider the Bourgeois contact structure coming from the following choice of open book on the sphere $\mathbb{S}^{2n-1}$. For $n\geq 3$, the open book coming from the $A_{1}-$singularity, with binding given by $B=\Sigma_{n-1}(2,2,\cdots,2)\cong S^{*}\mathbb{S}^{n-1}$; this is just the positive stabilisation $\mathbb{S}^{2n-1}=OB(D^{*}\mathbb{S}^{n-1},\tau)$ of the standard open book $OB(\mathbb{D}^{2n-2},\operatorname{Id})$, where $\tau$ is the Dehn-Seidel twist. For $n=2$, an open book $OB(\Sigma,\psi)$ for the standard contact structure on $\mathbb{S}^{3}$, whose page $\Sigma$ is non-sporadic and has homology $H_{1}(\Sigma)$ of rank at least $2$; for example, one can take repeated positive stabilisations of the trivial open book.
@@
Notice that, topologically, the Bourgeois manifold associated to these open books is of the form $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$. We then perform subcritical/flexible surgeries to topologically obtain a sphere $\mathbb{S}^{2n+1}$. More precisely, first observe that, by the $h$-principle for isotropic embeddings [EM02, Section 12.4] and for loose Legendrians [Mur12], one can always realise smooth surgeries of index $1$ and $2$ via contact surgeries. In view of this, we first perform two index $1$ surgeries that kill the generators of the fundamental group in the torus factor, to obtain a simply connected manifold with the homology of $\mathbb{S}^{2n-1}\times\mathbb{S}^{2}$. Notice that, since these generators are non-trivial in homology, the resulting cobordism will have trivial $c_{1}$. An index $2$ surgery then kills the generator in degree $2$ (which is indeed spherical by the Hurewicz theorem). In order to ensure that this surgery is possible one needs to make sure that the normal bundle is trivial: however, this is ensured by the assumption that $c_{1}$ vanishes, since the obstruction to the normal bundle being trivial is given by the second Stiefel-Whitney class $w_{2}$, to which $c_{1}$ reduces modulo $2$. Finally, the corresponding handle attachments applied to the boundary of the null-cobordism $\mathbb{D}^{2n}\times\mathbb{\mathbb{T}}^{2}$ yield a simply connected homology ball, which is then smoothly a ball by the $h$-cobordism theorem, giving that the manifold obtained by these surgeries is indeed a smoothly standard sphere. Realising these surgeries as Weinstein handle attachment according to the above mentioned $h$-principles, we thus obtain the desired contact structure $\xi_{ex}$ on $\mathbb{S}^{2n+1}$.
@@
Now, we claim that $(\mathbb{S}^{2n+1},\xi_{ex})$ is algebraically tight. Indeed, the Bourgeois structures before surgeries are algebraically tight: for $n=2$ this is the content of Corollary 3.6, while for $n\geq 3$ these structures are in fact $1$-ADC by Example 2.10, hence algebraically tight in view of Proposition 3.2. Then, since algebraic tightness is preserved under contact surgery, the claim follows.
Notice also that the almost contact structure underlying $\xi_{ex}$ is the standard one on $\mathbb{S}^{2n+1}$, since it extends as an almost complex structure on the smooth filling which is just a ball, on which any two almost complex structures are homotopic.
@@
To obstruct fillability, we want to apply Theorem 4.1 to the filling given by stacking the hypothetical filling of the contact structure on the sphere with the strong cobordism of Proposition 2.13. First, we remark that the $1$- surgeries can all performed along curves lying in a small neighbourhood of a fixed $\mathbb{T}^{2}$-fiber, since the isotropic $h$-principle is $C^{0}$-small and the $2$-surgery can be performed on a sphere contained in the trace of of this small neighbourhood under the initial surgery. In particular, all can be done on one side of the $\mathbb{S}^{1}$-equivariant decomposition for the Bourgeois contact structure. Now, we have $V_{+}=\Sigma\times D^{*}\mathbb{S}^{1}$ and $M^{\prime}_{+}$ is the resulting flexibly fillable contact manifold given by first capping then doing surgery as in Proposition 2.13. If $(\mathbb{S}^{2n+1},\xi_{ex})$ has a strong filling $W_{0}$, we get a strong filling $W$ of $M^{\prime}_{+}$ by stacking the strong cobordism of Proposition 2.13 on $W_{0}$. Note that there exists a homology class $\beta\in H_{n}(M^{\prime}_{+};\mathbb{Q})$ that is non-trivial in the flexible filling $W^{\prime}$ and is contained in $V_{+}\times\{pt\}$. Since we can always find a non-trivial homology class $\beta^{\prime}\in H_{n}(W^{\prime};\mathbb{Q})$ contained in $V_{+}\times\{pt\}$ by our assumption on $\Sigma$, we can take $\beta$ to be any preimage of $\beta$ under the surjective map $H_{n}(M^{\prime}_{+};\mathbb{Q})\to H_{n}(W^{\prime};\mathbb{Q})$. Since $V_{+}\times\{pt\}$ can be isotoped to $\mathbb{S}^{2n+1}$ in the cobordism, $\beta$ must be trivial in the glued strong filling contradicting Theorem 4.1 for $M=\emptyset$. ∎
@@
For concreteness we chose to take the open book determined by the $A_{1}$-singularity in the proof above. This said, any nontrivial Milnor open book whose binding is $-1$-ADC would suffice.
@@
Let $(\mathbb{S}^{2n+1},\xi_{ex})$ be a exotic contact sphere in Theorem B, then $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ is not strongly fillable for any contact manifold $M$ if $n\geq 3$. When $n=2$, the same holds as long as $c_{1}(M)$ is zero.
@@
Note that from the proof of Theorem 5.1, there exists a strong cobordism from $(\mathbb{S}^{2n+1},\xi_{ex})$ to a flexibly fillable contact manifold $M^{\prime}_{+}$ with vanishing first Chern class, such that there is a non-trivial homology class $\beta\in H_{n}(M^{\prime}_{+};\mathbb{Q})$ annihilated in the cobordism but not in the flexible filling. As the cobordism is obtained by blowing-down a Giroux domain and contact surgeries, which are operations performed away from a point, we get a strong cobordism from $(\mathbb{S}^{2n+1},\xi_{ex})\#M$ to $M^{\prime}_{+}\#M$ and the homology class corresponding to $\beta$ is again annihilated in the cobordism. We then apply Theorem 4.1 to obstruct fillability by stacking the blow-up cobordism on top of any hypothetical filling of $(\mathbb{S}^{2n+1},\xi_{ex})\#M$. ∎
@@
We can assume $(M,\xi)$ admits a strong filling, for otherwise, we have nothing to prove. In this case the connected sum $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ is not strongly fillable by Theorem 5.3. Finally by Proposition 3.2, $(M,\xi)$, hence $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ are both algebraically tight. ∎
@@ 5.2. Abundance of exotic structures on spheres
When $n\geq 3$, the contact manifold $(\mathbb{S}^{2n+1},\xi_{ex})$ is $(n-3)$-ADC.
@@
Lemma 2.8 gives the indices of contractible Reeb orbits on the Bourgeois manifold $BO(D^{*}\mathbb{S}^{n-1},\tau)$. On the other hand, the subcritial surgeries kill part of the topology of $BO(D^{*}\mathbb{S}^{n-1},\tau)$ making non-contractible orbits of $BO(D^{*}\mathbb{S}^{n-1},\tau)$ contractible in $(\mathbb{S}^{2n+1},\xi_{ex})$. We then need to compute the Conley-Zehnder indices of them as well. For simplicity we pick a global trivialisation of the almost contact structure underlying the Bourgeois contact structure, which is split as an almost contact structure according to Remark 2.2, by simply taking a trivialisation of $(\mathbb{S}^{2n-1},\xi_{st})$ and trivialising the tangent bundles of the $\mathbb{T}^{2}$-fibers in an equivariant manner. The Reeb orbits on $BO(D^{*}\mathbb{S}^{n-1},\tau)$ then have Conley-Zehnder indices as follows:
@@
    On $B\times D^{*}\mathbb{T}^{2}$, where $B=S^{*}\mathbb{S}^{n-1}$, the computation of the Conley-Zehnder indices is exactly the same as in Lemma 2.8. Namely, we have
    $\mu_{CZ}(\gamma_{q})=\mu_{CZ}(\gamma_{B})-2+\mbox{ind}_{q}(H)\geq\mu_{CZ}(\gamma_{B}).$
    Here, $\gamma_{q}=\gamma_{B}\times\{(q,r=0)\}$ is the contractible orbit on the binding corresponding to the critical point $q\in\mathbb{T}^{2}\times\{0\}$ of a Morse function $H$ on $D^{*}\mathbb{T}^{2}$. From the computations in [KvK16, p. 37] for the $A_{1}$-singularity, we obtain that $\mu_{CZ}(\gamma_{B})\geq n-2$, and thus
    $\mu_{CZ}(\gamma_{q})\geq n-2,$
@@
    On the product $\mbox{Map}(D^{*}\mathbb{S}^{n-1},\tau)\times\mathbb{T}^{2}$, the orbits stay tangent to the pages of the open book. Moreover, for a fixed angle $\theta\in\mathbb{S}^{1}$, the closed Reeb orbits tangent to the $\theta-$page are naturally parametrized by the page $D^{*}\mathbb{S}^{n-1}$ itself, namely they are flat geodesics on $\mathbb{T}^{2}$ with rational slope $\theta$ (recall Formula (2)), and so they form an $\mathbb{S}^{1}$ Morse–Bott family when viewed as Reeb orbits on $\mathbb{T}^{3}$. As in Lemma 2.8, we locally perturb the contact form by a Morse function with a unique local maximum on $D^{*}\mathbb{S}^{n-1}$ and two critical points $q_{1},q_{2}$ lying in the zero section, respectively of index $22n-$ 2 and $n-1$. The minimal Conley-Zehnder index of the orbits on $\mathbb{T}^{3}$ after perturbing with a Morse function on $\mathbb{S}^{1}$ with critical points $p_{1},p_{2}$ is zero. By standard properties of the Conley-Zehnder index, if we denote by $\gamma_{p_{j}}$ the Reeb orbit on $\mathbb{T}^{3}$ corresponding to $p_{j}$, then the non-degenerate orbit $\gamma_{q_{i},p_{j}}$ corresponding to the pair $(q_{i},p_{j})$ has Conley-Zehnder index
@@
    Therefore the Conley-Zehnder index of every orbit is non-negative along this region, and their SFT degrees are at least $n-2$.
@@
    Along the interface region $S^{*}\mathbb{S}^{n-1}\times\mathbb{T}^{3}$, the situation is modelled as a smoothened corner of the product $D^{*}\mathbb{S}^{n-1}\times D^{*}\mathbb{T}^{2}$. Computations in this setting are carried out in detail in [Zho21b, Proposition 6.18 (3)], and the result is that the minimal Conley-Zehnder index is the sum of minimal Conley-Zehnder indices on $S^{*}\mathbb{S}^{n-1}$ (i.e. $n-2$) and $S^{*}\mathbb{T}^{2}$ (i.e. 00) [^3] Note that [Zho21b] considered the $S^{1}$-family of Reeb orbits, whose generalized Conley-Zehnder index has an extra $1/2$. But here we consider non-degenerate Reeb orbits after a small perturbation of the Morse-Bott family, hence the minimal Conley-Zehnder index adds.. Hence we have $\mu_{CZ}(\gamma)\geq n-2$ for every orbit $\gamma$ in this region.
@@
    The attaching of two $1$-handles and a $2$-handles create Reeb orbits of index at least $n-2$. When we attach the two $1$-handles to kill the fundamental group of $BO(D^{*}\mathbb{S}^{n-1},\tau)$, to maintain the correctness of Conley-Zehnder indices computed above, we need to choose the framing so that the induced trivialization of the contact structure (used in cases 2 and 3 above) along the isotropic circle extends to a trivialization of the core disk in the handle. In fact since the $\mathbb{T}^{2}$-fibers over the binding of the open book in the base are isotropic, these give natural choices of isotropic circles for surgery (of any slope). One then simply takes the framing of the normal bundle induced from the trivialisation on the symplectic normal induced from the projection to the base.
In total we see that after surgery all periodic orbits have degree at least $n+1-3=n-2$ (up to a certain action threshold). Thus the resulting contact structure is $(n-3)$-ADC. ∎
As proved in [Laz20a], the ADC condition is sufficient to obtain well-defined invariants of Liouville-fillable contact manifolds using positive symplectic cohomology (generated by contractible orbits) of any of its topologically simple [^4] That is $c_{1}=0$ and the inclusion of the boundary induces an injective map on $\pi_{1}$. Liouville fillings. Under a slightly stronger assumption, namely that there is a non-degenerate Reeb vector field all of whose contractible orbits have SFT degree at least $2$, Cieliebak-Oancea [CO18, Section 9.5] described a version of symplectic homology on the trivial symplectic cobordism $M\times[0,1]$ associated to the contact manifold $M$. This yields a contact invariant in view of a neck-stretching argument. In fact, for exactly the same reason as in the fillable case [Laz20a], it is enough to have $1$-ADC for this to be a well-defined invariant of the contact manifold itself. With a slight abuse of notation, we will then denote the positive symplectic cohomology of a $1$-ADC contact manifold $(M,\xi)$ simply by $SH_{+}^{*}(M)$ (i.e. $SH^{>0}_{*}(M)$ in [CO18, Section 9.5]). When $M$ is the contact boundary of a topologically simple Liouville filling $W$, we have $SH_{+}^{*}(M)=SH_{+}^{*}(W)$.

### 14 Mar 2026 v5

@@ 5.1. Tight non-fillable structures on spheres
We first construct tight, non-fillable contact structures on $\mathbb{S}^{2n+1}$, for $n\geq 2$ to prove the following stronger version of Theorem B.
@@
Note that Theorem 5.1 directly gives Theorem B because of the general fact that algebraic tightness implies tightness, as recalled in Proposition 3.2 (1).
@@
Step 1: the geometric construction. We consider the Bourgeois contact structure coming from the following choice of open book on the sphere $\mathbb{S}^{2n-1}$. For $n\geq 3$, we chose the open book coming from the $A_{1}$-singularity, with binding given by $B=\Sigma_{n-1}(2,2,\cdots,2)\cong S^{*}\mathbb{S}^{n-1}$; this is just the positive stabilization $\mathbb{S}^{2n-1}=OB(D^{*}\mathbb{S}^{n-1},\tau)$ of the standard open book $OB(\mathbb{D}^{2n-2},\operatorname{Id})$, where $\tau$ is the Dehn-Seidel twist, as defined in detail in [Sei03] following the idea in [Arn95]. For $n=2$, we chose an open book $OB(\Sigma,\psi)$ for the standard contact structure on $\mathbb{S}^{3}$, whose page $\Sigma$ is non-sporadic and has homology $H_{1}(\Sigma)$ of rank at least $2$; for example, one can take repeated positive stabilizations of the trivial open book.
@@
Notice that, topologically, the Bourgeois manifold associated to these open books is of the form $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$. We then perform subcritical/flexible surgeries to topologically obtain a sphere $\mathbb{S}^{2n+1}$. More precisely, first observe that, by the $h$-principle for isotropic embeddings [EM02, Section 12.4] and for loose Legendrians [Mur12] (see also [CE12, Theorem 7.16]), one can always realize smooth surgeries of index $1$ and $2$ via contact surgeries. In view of this, we first perform two index $1$ surgeries that kill the generators of the fundamental group in the torus factor, to obtain a simply connected manifold with the homology of $\mathbb{S}^{2n-1}\times\mathbb{S}^{2}$. Notice that, since these generators are non-trivial in homology and we start from a Bourgeois contact structure with trivial $c_{1}$, the resulting cobordism will have trivial $c_{1}$.
Now, observe that the generator of the second homology group of the obtained manifold is spherical by the Hurewicz theorem. Moreover, using the fact that the ambient dimension is at least $5$, we can perturb a continuous map $f\colon\mathbb{S}^{2}\to M$ representing such a generator to a smooth embedding $\widetilde{f}\colon\mathbb{S}^{2}\to M$ according to standard transversality theory (see e.g. [Hir94, Theorems 2.6 and 2.13]). We can then perform an index $2$ surgery along $\widetilde{f}(\mathbb{S}^{2})$, which has the effect of killing the generator in degree-$2$ homology.
In order to ensure that this surgery is possible, one needs to make sure that the normal bundle is trivial; however, this is ensured by the assumption that $c_{1}$ vanishes, since the obstruction to the normal bundle being trivial is given by the second Stiefel-Whitney class $w_{2}$, to which $c_{1}$ reduces modulo $2$. Finally, the corresponding handle attachments applied to the boundary of the null-cobordism $\mathbb{D}^{2n}\times\mathbb{\mathbb{T}}^{2}$ yield a simply connected homology ball $X$, which is then smoothly a ball by the $h$-cobordism theorem, giving that the manifold obtained by these surgeries is indeed a smoothly standard sphere. Realizing these surgeries as Weinstein handle attachment according to the above-mentioned $h$-principles, we thus obtain the desired contact structure $\xi_{ex}$ on $\mathbb{S}^{2n+1}$.
@@
Step 2: algebraic tightness. Now, we claim that $(\mathbb{S}^{2n+1},\xi_{ex})$ is algebraically tight. Indeed, the Bourgeois structures before surgeries are algebraically tight: for $n=2$ this is the content of Corollary 3.6, while for $n\geq 3$ these structures are in fact $1$-ADC by Example 2.10, hence algebraically tight in view of Proposition 3.2. Then, since algebraic tightness is preserved under contact surgery, the claim follows.
Step 3: the almost contact class. As explained in Step 1, there is a Weinstein cobordism from a Bourgeois structure on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$, associated to the standard contact structure on $\mathbb{S}^{2n-1}$ (for some open book that is here irrelevant), to the contact sphere $(\mathbb{S}^{2n+1},\xi_{ex})$. According to Remark 2.2, this is in particular an almost complex cobordism from a split almost contact structure $\eta:=(\xi_{std},d\alpha_{std})\oplus(T\mathbb{T}^{2},\Omega)$ to the underlying almost contact structure $\zeta$ to $\xi_{ex}$ on $\mathbb{S}^{2n+1}$. Now, as $\eta$ admits as almost complex filling $(B^{2n}\times\mathbb{T}^{2},\omega_{std}\oplus\Omega)$, this gives an almost complex filling $(X,J)$ of $(\mathbb{S}^{2n+1},\xi_{ex})$, where $X$ is exactly the simply connected homology ball, and hence smoothly a ball by the h-cobordism theorem, that was described in Step 1. In particular, as there is only one almost complex structure up to homotopy on the ball, the almost contact structure $\zeta$ underlying $\xi_{ex}$ is homotopic to the standard one on $\mathbb{S}^{2n+1}$, as desired.
@@
Step 4: non-fillability. To obstruct fillability, we want to apply Theorem 4.1 to the filling given by stacking the hypothetical filling of the contact structure on the sphere with the strong cobordism of Proposition 2.13, and so we first check the hypotheses of this proposition. First, we remark that the $1$-surgeries can all be done along curves lying in a small neighborhood of a fixed $\mathbb{T}^{2}$-fiber, since the isotropic $h$-principle is $C^{0}$-small, and that the $2$-surgery can be performed on a sphere contained in the trace of this small neighborhood under the initial surgery. In particular, all can be done on one side of the $\mathbb{S}^{1}$-equivariant decomposition for the Bourgeois contact structure. Now, we have $V_{+}=\Sigma\times D^{*}\mathbb{S}^{1}$ and $M^{\prime}_{+}$ is the resulting flexibly fillable contact manifold given by first capping then doing surgery as in Proposition 2.13. The flexible filling $W_{0}$ of $M^{\prime}_{+}$ is obtained from attaching two $2$-handles and one flexible $3$-handle along a loose Legendrian sphere to the subcritical domain $V_{+}\times\mathbb{D}^{2}$. If $(\mathbb{S}^{2n+1},\xi_{ex})$ has a strong filling $W$, we get a strong filling $W^{\prime}$ of $M^{\prime}_{+}$ by stacking the strong cobordism of Proposition 2.13 on $W$. Note that there exists a homology class $\beta\in H_{n}(M^{\prime}_{+};\mathbb{Q})$ that is non-trivial in the flexible filling $W_{0}$ and is contained in $V_{+}\times\{pt\}$. Indeed, since we can always find a non-trivial homology class $\beta^{\prime}\in H_{n}(W_{0};\mathbb{Q})$ contained in $V_{+}\times\{pt\}$ by our assumption on $\Sigma$ (recall that we arranged, at the beginning of the proof, that $H_{1}(\Sigma)$ has rank at least $2$ when $n=2$), we can take $\beta$ to be any preimage of $\beta^{\prime}$ under the surjective map $H_{n}(M^{\prime}_{+};\mathbb{Q})\to H_{n}(W_{0};\mathbb{Q})$. Since $V_{+}\times\{pt\}$ can be isotoped to $\mathbb{S}^{2n+1}$ in the cobordism, $\beta$ must be trivial in the glued strong filling contradicting Theorem 4.1 for $M=\emptyset$. ∎
@@
For concreteness, we chose to take the open book determined by the $A_{1}$-singularity in the proof above. This said, any nontrivial open book whose binding is (1)(-1)-index positive and such that the page has a non-trivial middle homology group would suffice.
@@
Let $(\mathbb{S}^{2n+1},\xi_{ex})$ be an exotic contact sphere in Theorem 5.1. Then $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ is not strongly fillable for any contact manifold $M$ if $n\geq 3$. When $n=2$, the same holds if $c_{1}(M)$ vanishes.
@@
Note that from the proof of Theorem 5.1 (Step 4), there exists a strong cobordism from $(\mathbb{S}^{2n+1},\xi_{ex})$ to a flexibly fillable contact manifold $M^{\prime}_{+}$ with vanishing first Chern class, such that there is a non-trivial homology class $\beta\in H_{n}(M^{\prime}_{+};\mathbb{Q})$ annihilated in the cobordism but not in the flexible filling. As the cobordism is obtained by blowing-down a Giroux domain and contact surgeries, which are operations performed away from a point, we get a strong cobordism from $(\mathbb{S}^{2n+1},\xi_{ex})\#M$ to $M^{\prime}_{+}\#M$ and the homology class corresponding to $\beta$ is again annihilated in the cobordism. We then apply Theorem 4.1 to obstruct fillability by stacking the blow-up cobordism on top of any hypothetical filling of $(\mathbb{S}^{2n+1},\xi_{ex})\#M$. (Note that the additional assumption of $c_{1}(M)=0$ in the case of $n=2$ is needed in order to apply Theorem 4.1.(a) in this non-subcritical case.) ∎
@@
We can assume $(M,\xi)$ admits a strong filling, for otherwise, we have nothing to prove. In this case the connected sum $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ is not strongly fillable by Theorem 5.3. We claim that algebraic tightness follows from the properties listed in Section 3.
Indeed, by Proposition 3.2 (4), we know that $(M,\xi)$ is algebraically tight. Then, by Proposition 3.2 (2) and Theorem 5.1, the disjoint union $(M,\xi)\sqcup(\mathbb{S}^{2n+1},\xi_{ex})$ is algebraically tight as well. Since there is a Weinstein cobordism from $(M,\xi)\sqcup(\mathbb{S}^{2n+1},\xi_{ex})$ to the connected sum $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ by a Weinstein $1$-handle, $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ is algebraically tight by Proposition 3.2 (3). ∎
@@ 5.2. Abundance of exotic structures on spheres
When $n\geq 3$, the contact manifold $(\mathbb{S}^{2n+1},\xi_{ex})$ is (n3)(n-3)-ADC.
@@
Lemma 2.8 gives the indices of contractible Reeb orbits on the Bourgeois manifold $BO(D^{*}\mathbb{S}^{n-1},\tau)$. On the other hand, the subcritical surgeries kill part of the topology of $BO(D^{*}\mathbb{S}^{n-1},\tau)$, making non-contractible orbits of $BO(D^{*}\mathbb{S}^{n-1},\tau)$ contractible in $(\mathbb{S}^{2n+1},\xi_{ex})$. We then need to compute the Conley-Zehnder indices of them as well.
For simplicity we pick a global trivialisation of the almost contact structure underlying the Bourgeois contact structure, which is split as an almost contact structure according to Remark 2.2, by simply taking a trivialization of $(\mathbb{S}^{2n-1},\xi_{st})$ and trivializing the tangent bundles of the $\mathbb{T}^{2}$-fibers in an equivariant manner. The Reeb orbits on $BO(D^{*}\mathbb{S}^{n-1},\tau)$ then have Conley-Zehnder indices as follows:
@@
    On $B\times D^{*}\mathbb{T}^{2}$, where $B=S^{*}\mathbb{S}^{n-1}$, the contact form is Morse-Bott. The computation of the Conley-Zehnder indices is in Lemma 2.8. Namely, we have
    $\mu_{LCZ}(\gamma_{q})=\mu_{LCZ}(\gamma_{B}).$
    Here, $\gamma_{q}=\gamma_{B}\times\{(r=0,q)\}$ is the contractible orbit on the binding. And $\gamma_{B}$ is the Reeb orbits on $B$, which comes in a Morse-Bott family. From the computations in [KvK16, p. 37] for the $A_{1}$-singularity, we obtain that $\mu_{LCZ}(\gamma_{B})\geq n-2$, and thus
    $\mu_{LCZ}(\gamma_{q})\geq n-2,$
@@
    On the product $\mbox{Map}(D^{*}\mathbb{S}^{n-1},\tau)\times\mathbb{T}^{2}$, we have the following situation. Recall from Theorem 2.1, that the Bourgeois form $\alpha_{BO}$ in this region simply can be written as $\pi^{*}\alpha_{T}+\lambda_{fib}$ on the total space of the natural projection $\pi\colon\mathrm{Map}(D^{*}\mathbb{S}^{n-1},\tau)\times\mathbb{T}^{2}\to\mathbb{S}^{1}\times\mathbb{T}^{2}=\mathbb{T}^{3}$, where $\alpha_{T}=Cd\theta+\cos\theta\,dx-\sin\theta\,dy$ on $\mathbb{T}^{3}\ni(\theta,x,y)$ and $\lambda_{fib}$ is a $1$-form that is Liouville on each fiber of $\pi$. Moreover, it follows from Equation (2) that in this region the Reeb vector field $R$ is just a lift via $d\pi$ of the Reeb vector field $R_{T}$ of $\alpha_{T}$ on $\mathbb{T}^{3}$. More precisely, if $(p,x,y)\in\mathrm{Map}(D^{*}\mathbb{S}^{n-1},\tau)\times\mathbb{T}^{2}$ lies in the fiber over $(\theta,x,y)\in\mathbb{T}^{3}$, the Reeb orbit through it is tangent to $\{p\}\times\mathbb{T}^{2}$ and is a geodesic for the flat metric, that has starting vector at angle (w.r.t. the coordinates $(x,y)$) exactly $\theta$. In other words, Reeb orbits are naturally organized in $D^{*}\mathbb{S}^{n-1}\times\mathbb{S}^{1}$-Morse–Bott families, each of which projects down via $\pi$ to a $\mathbb{S}^{1}$-Morse–Bott family of orbits of $R_{T}$ on $\mathbb{T}^{3}$.
    Now, perturbing with a Morse function on $\mathbb{S}^{1}$ with critical points $p_{1},p_{2}$, one can find another contact form $\alpha_{T}^{\prime}$ that is $C^{\infty}$-close to $\alpha_{T}$ on $\mathbb{T}^{3}$ such that each $\mathbb{S}^{1}$-Morse–Bott family of Reeb orbits corresponds to a pair of Reeb orbits. We denote by $\gamma_{p_{j}}$ the Reeb orbit on $\mathbb{T}^{3}$ corresponding to $p_{j}$. By standard properties of the Conley-Zehnder index, the minimal Conley-Zehnder index of the orbits on $\mathbb{T}^{3}$ after such perturbation is zero.
    We can then look at the perturbation $\alpha_{BO}^{\prime}$ given by $\pi^{*}\alpha_{T}^{\prime}+\lambda_{fib}$. Likewise, we can further locally perturb the contact form by a Morse function with a unique local maximum on $D^{*}\mathbb{S}^{n-1}$ and two critical points $q_{1},q_{2}$ lying in the zero section, respectively of index 2n22n-2 and n1n-1. We denote by $\gamma_{q_{i},p_{j}}$ the non-degenerate orbit corresponding to the pair $(q_{i},p_{j})$. Again, by standard properties of the Conley-Zehnder index, we then have:
@@
    Therefore, the Conley-Zehnder index of every orbit is non-negative along this region $\mathrm{Map}(D^{*}\mathbb{S}^{n-1},\tau)\times\mathbb{T}^{2}$, and their SFT degrees are at least n2n-2.
@@
    Along the interface region $S^{*}\mathbb{S}^{n-1}\times\mathbb{T}^{3}$, the situation is modeled as a smoothened corner of the product $D^{*}\mathbb{S}^{n-1}\times D^{*}\mathbb{T}^{2}$. Computations in this setting are carried out in detail in [Zho21b, Proposition 6.18 (3)], and the result is that the minimal Conley-Zehnder index is the sum of minimal Conley-Zehnder indices on $S^{*}\mathbb{S}^{n-1}$ (i.e. n2n-2) and $S^{*}\mathbb{T}^{2}$ (i.e. 0). (Note that [Zho21b] considered the $S^{1}$-family of Reeb orbits, whose generalized Conley-Zehnder index has an extra $1/2$. On the other hand, here we consider non-degenerate Reeb orbits after a small perturbation of the Morse-Bott family, hence the minimal Conley-Zehnder index adds.) Hence, we have $\mu_{CZ}(\gamma)\geq n-2$ for every orbit $\gamma$ in this region.
@@
    The attaching of two $2$-handles and a $3$-handles create contractible new Reeb orbits, but these have Conley-Zehnder indices at least n1n-1 asymptotically, by the proof of [Laz20a, Theorem 3.14] (or [Zho23c, Proposition 2.3] for more precise numbers). When we attach the two $2$-handles to kill the fundamental group of $BO(D^{*}\mathbb{S}^{n-1},\tau)$. Here, to maintain the correctness of Conley-Zehnder indices computed above, we need to choose the framing of the symplectic normal bundle of the isotropic circles so that the induced trivialization of the contact structure (used in cases 2 and 3 above) along the isotropic circle extends to a trivialization of the core disk in the handle. In fact, since the $\mathbb{T}^{2}$-fibers over the binding of the open book in the base are isotropic, these give natural choices of isotropic circles for surgery (of any slope). One then simply takes the framing of the normal bundle induced from the trivialisation of the symplectic normal bundle induced from the projection to the base.
In total, we have shown that the Bourgeois contact form considered in §2 is (n3)(n-3)-index positive even including all the non-contractible Reeb orbits with a suitable framing (called $\pi_{1}$-sensitive ADC in §6). In particular, it is (n3)(n-3)-ADC after a small perturbation as in [Bou02b, Lemma 2.3]. The framing on the Bourgeois contact structure is compatible with framing on $\mathbb{S}^{2n+1}$ after the surgery as explained in point (4), the subcritical surgery will preserve the (n3)(n-3)-ADC property by [Laz20a, Theorems 3.14] [^7] As the proof of [Laz20a, Theorem 3.14] shows that index $1$ or $2$ subcritical handle to a $(2n+1)$-dimensional contact manifold adds Reeb orbits of Conley-Zehnder index at least $n$ asymptotically. It would preserve $p$-ADC as long as p<2n2p<2n-2., or Proposition 6.4 below. ∎
As proved in [Laz20a], the ADC condition is sufficient to obtain well-defined invariants of Liouville-fillable contact manifolds using positive symplectic cohomology (generated by contractible orbits) of any of its topologically simple Liouville fillings (that is, Liouville fillings with $c_{1}=0$ and so that the inclusion of the boundary induces an injective map on $\pi_{1}$). Under a slightly stronger assumption, namely that there is a non-degenerate Reeb vector field all of whose contractible orbits have SFT degree at least $2$, Cieliebak-Oancea [CO18, Section 9.5] described a version of symplectic homology on the trivial symplectic cobordism $M\times[0,1]$ associated to the contact manifold $M$. The latter construction yields a contact invariant in view of a neck-stretching argument. In fact, for exactly the same reason as in the fillable case [Laz20a], it is enough to have $1$-ADC for this to be a well-defined invariant of the contact manifold itself. With a slight abuse of notation, we will then denote the positive symplectic cohomology of a $1$-ADC contact manifold $(M,\xi)$ simply by $SH_{+}^{*}(M)$ (i.e. $SH^{>0}_{*}(M)$ in [CO18, Section 9.5]). When $M$ is the contact boundary of a topologically simple Liouville filling $W$, we have $SH_{+}^{*}(M)=SH_{+}^{*}(W)$.